x^3y''' + 2x^2y'' = x + sin(lnx)


1
Expert's answer
2022-01-19T08:18:51-0500

"y''=v,y'''=v'"

"v'+2v\/x=(x+sin(lnx))\/x^3"


"\mu(x)=e^{\int 2/x dx}=x^2"

multiply both sides by "\mu(x)" :

"x^2v'+2xv=(x+sin(lnx))\/x"


"\frac{d}{dx}(x^2v)=(x+sin(lnx))/x"


"x^2v=\int(x+sin(lnx))/xdx"


"x^2v=x-cos(lnx)+c_1"

"v=y''=(x-cos(lnx)+c_1)\/x^2"


"y'=\int(x-cos(lnx)+c_1)/x^2dx"


"\int \frac{cos(lnx)}{x^2}dx=|u=lnx,dx=xdu|=\int e^{-u}cosudu="


"=-e^{-u}cosu-\intop e^{-u}cosudu=-e^{-u}cosu+-e^{-u}sinu-\intop e^{-u}cosudu"


"\intop e^{-u}cosudu=\frac{cos(lnu)-sin(lnxu)}{2}=\frac{cos(lnx)-sin(lnx)}{2x}"


"y'=\frac{cos(lnx)-sin(lnx)}{2x}+lnx-c_1/x+c_2"


"y=\int (\frac{cos(lnx)-sin(lnx)}{2x}+lnx-c_1/x+c_2)dx"


"\int \frac{cos(lnx)-sin(lnx)}{x}dx=\int (cos(lnx)-sin(lnx))d(lnx)=sin(lnx)+cos(lnx)"


"y=\frac{cos(lnx)+sin(lnx)}{2}-x+xlnx-c_1lnx+c_2x+c_3"